K’s Fashions is growing quickly. Dividends are expected to increase by 8 percent annually for the next three years, with the growth rate falling off to a constant 3 percent thereafter. The required return is 14 percent and the company just paid a $3.64 annual dividend. What is the current share price?
Answer #1

D1=(3.64*1.08)=$3.9312

D2=(3.9312*1.08)=$4.245696

D3=(4.245696*1.08)=$4.58535168

Value after year 3=(D3*Growth Rate)/(Required Rate-Growth rate)

=(4.58535168*1.03)/(0.14-0.03)

=$42.93556573

Hence current share price=Future dividends*Present value of discounting factor(rate%,time period)

=$3.9312/1.14+$4.245696/1.14^2+$4.58535168/1.14^3+$42.93556573/1.14^3

which is equal to

=$38.79(Approx).